# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

### v1

@@ Appendix A A Tensor Product of Virasoro Algebra Representations
In this appendix we show the solutions of equation (2.8) and (2.9e).
@@
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=1,$ $l_{k}=l_{i}+l_{j},$ (2.9ca)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{k}=l_{i}+l_{j}-1,$ (2.9cb)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(l_{j}r_{i}+r_{j}),$ $l_{i}=0,\quad l_{k}=l_{j}-2,$ (2.9cc)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=(r_{i}+r_{j})(l_{j}r_{i}-l_{i}r_{j}),$ $l_{i}+l_{j}=1,\quad l_{k}=-1,$ (2.9cd)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9ce)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{i}=l_{j}=0,\quad l_{k}=-3,$ (2.9cf)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(r_{i}+r_{j})(2r_{i}+r_{j}),$ $l_{i}=0,\quad l_{j}=2,\quad l_{k}=-1.$ (2.9cg)
@@
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}+r_{j}},$ $l_{i}+l_{j}=-1,\quad l_{k}=0,$ (2.9da)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}},$ $l_{i}=-1,\quad l_{j}=l_{k},$ (2.9db)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta(r_{i}+r_{j}),$ $l_{i}=l_{j}=l_{k}=0,$ (2.9dc)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta_{r_{j}},$ $l_{i}=0,\quad l_{k}=l_{j}=-1,$ (2.9dd)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0,$ (2.9de)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\frac{1}{r_{i}}\delta_{r_{i}+r_{j}}+\frac{1}{r_{j}}\delta_{r_{i}}-\frac{1}{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0.$ (2.9df)
These solutions remain valid under the exchange of $i$ and $j$. Among them, (2.9ca), (2.9cb), and (2.9cc) coincide with the solutions obtained in (2.9). Solution (2.9df) does not play a role in the construction of the superalgebras. And the other solutions are only consistent for negative parameter values, so they are not in our interest for constructing the superalgebras.
It is worth emphasizing that (2.9da) is the solution used in (2.9h), which automatically satisfied by the canonical commutation relations of the field modes. For instance, (2.9i) requires $l_{\varphi}+l_{\phi}=-1$ and $\bar{l}_{\varphi}+\bar{l}_{\phi}=-1$, while the spin indices of the identity operator are $l_{I}=\bar{l}_{I}=0$.
The difference equation for the field modes (2.9d) is repeated here:
$((l_{i}+1)m-r)c(r-m,t)+(l_{j}m-t+r)c(r,t)-(l_{Q}m-t)c(r,t+m)=0.$ (2.9e)
@@
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r},$ $l_{i}=0,\quad l_{k}=l_{j},$ (2.9ha)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=n\delta_{r},$ $l_{i}=l_{k}=0,\quad l_{j}=-1,$ (2.9hb)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\frac{1}{n}\delta_{n-r}-\frac{1}{r}\delta_{n}-\frac{1}{n}\delta_{r},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9hc)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r}\delta_{n},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (2.9hd)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{n},$ $l_{i}+l_{j}=l_{k}=-1,$ (2.9he)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=r\delta_{n},$ $l_{i}=l_{j}=l_{k}=-1.$ (2.9hf)
These solutions remain valid under the exchange $l_{i}\leftrightarrow l_{j}$ accompanied by the shift $n\to n-r$. Among them, only (2.9ga) and (2.9gb), which correspond to (2.9e), are relevant for constructing the field modes.

### v2

@@ Appendix A A Tensor Product of Virasoro Algebra Representations
In this appendix we show the solutions of equation (2.8) and (3.5).
@@
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=1,$ $l_{k}=l_{i}+l_{j},$ (A.3a)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{k}=l_{i}+l_{j}-1,$ (A.3b)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(l_{j}r_{i}+r_{j}),$ $l_{i}=0,\quad l_{k}=l_{j}-2,$ (A.3c)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=(r_{i}+r_{j})(l_{j}r_{i}-l_{i}r_{j}),$ $l_{i}+l_{j}=1,\quad l_{k}=-1,$ (A.3d)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (A.3e)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=l_{j}r_{i}-l_{i}r_{j},$ $l_{i}=l_{j}=0,\quad l_{k}=-3,$ (A.3f)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}(r_{i}+r_{j})(2r_{i}+r_{j}),$ $l_{i}=0,\quad l_{j}=2,\quad l_{k}=-1.$ (A.3g)
@@
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}+r_{j}},$ $l_{i}+l_{j}=-1,\quad l_{k}=0,$ (A.4a)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}},$ $l_{i}=-1,\quad l_{j}=l_{k},$ (A.4b)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta(r_{i}+r_{j}),$ $l_{i}=l_{j}=l_{k}=0,$ (A.4c)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=r_{i}\delta_{r_{j}},$ $l_{i}=0,\quad l_{k}=l_{j}=-1,$ (A.4d)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\delta_{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0,$ (A.4e)
$f_{l_{i},l_{j}}^{l_{k}}(r_{i},r_{j})=\frac{1}{r_{i}}\delta_{r_{i}+r_{j}}+\frac{1}{r_{j}}\delta_{r_{i}}-\frac{1}{r_{i}}\delta_{r_{j}},$ $l_{i}=l_{j}=-1,\quad l_{k}=0.$ (A.4f)
These solutions remain valid under the exchange of $i$ and $j$. Among them, (A.3a), (A.3b), and (A.3c) coincide with the solutions obtained in (2.9). Solution (A.4f) does not play a role in the construction of the superalgebras. And the other solutions are only consistent for negative parameter values, so they are not in our interest for constructing the superalgebras.
It is worth emphasizing that (A.4a) is the solution used in (3.8), which automatically satisfied by the canonical commutation relations of the field modes. For instance, (3.9) requires $l_{\varphi}+l_{\phi}=-1$ and $\bar{l}_{\varphi}+\bar{l}_{\phi}=-1$, while the spin indices of the identity operator are $l_{I}=\bar{l}_{I}=0$.
The difference equation for the field modes (3.4) is repeated here:
$((l_{i}+1)m-r)c(r-m,t)+(l_{j}m-t+r)c(r,t)-(l_{Q}m-t)c(r,t+m)=0.$ (A.5)
@@
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r},$ $l_{i}=0,\quad l_{k}=l_{j},$ (A.8a)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=n\delta_{r},$ $l_{i}=l_{k}=0,\quad l_{j}=-1,$ (A.8b)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\frac{1}{n}\delta_{n-r}-\frac{1}{r}\delta_{n}-\frac{1}{n}\delta_{r},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (A.8c)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{r}\delta_{n},$ $l_{i}=l_{j}=0,\quad l_{k}=-1,$ (A.8d)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=\delta_{n},$ $l_{i}+l_{j}=l_{k}=-1,$ (A.8e)
$c_{l_{i},l_{j}}^{l_{k}}(r,n)=r\delta_{n},$ $l_{i}=l_{j}=l_{k}=-1.$ (A.8f)
These solutions remain valid under the exchange $l_{i}\leftrightarrow l_{j}$ accompanied by the shift $n\to n-r$. Among them, only (A.7a) and (A.7b), which correspond to (3.5), are relevant for constructing the field modes.